# What Comes After Harm? Mapping Reparative Actions in AI through Justice Frameworks

### v1

@@ Acknowledgment
Advocacy groups frequently turn to litigation when other accountability channels fail. For instance, the UK-based Public Law Project challenged the Home Office’s algorithm for flagging “sham marriages,” citing discrimination and GDPR violations. As of the latest reporting, the case had not reached a resolution (A1389).
Affected Individuals, including public figures and professionals, also initiated lawsuits, typically over severe harm or personal rights violations. For example, several religious authors sued Meta and Microsoft for allegedly generating AI content derived from their copyrighted works (A1149). In another case, Megan Garcia filed a negligence lawsuit against Character AI after her son died by suicide following emotional attachment to a chatbot (A1781). Actress Scarlett Johansson also sued an AI developer for unauthorized use of her likeness (A1165).
@@
While lawsuits provide a formal avenue for redress, our findings show they are typically slow-moving, legally complex, and more accessible to well-resourced individuals or organizations. Cases initiated by regulators are far more likely to result in concrete consequences than those led by private actors.
@@
Removal of content and product recall
@@
While our taxonomy identifies which stakeholder groups initiate each type of reparative action, it is organized around actions and does not reveal how individual stakeholders contribute across the broader set of actions. To complement this, Table 2 presents the frequency with which each stakeholder group initiates each action. This comparative view highlights patterns of responsibility and exposes disparities in who drives different forms of reparation.
Stakeholder Group Outcry Comm. Invest. Lawsuit Repr. Comp. Design Recall Discont. Policy
Perpetrators – AI Corporations 0 46.1% 6.8% 0.9% 1.1% 5.0% 10.3% 9.2% 3.7% 2.7%
Perpetrators – End-Users 0.1% 7.2% 0.7% 0.4% 0.3% 0.9% 0.1% 1.5% 0.5% 0.4%
Affected Users 2.6% 0 0.1% 7.9% 0 0.1% 0 0 0 0
Affected Non-Users 0.6% 0 0 2.6% 0 0 0 0 0 0
General Public 6.6% 0 0 0 0 0 0 0 0 0
Regulators and Government Bodies 0.2% 0.3% 20.8% 2.2% 9.5% 1.2% 0.8% 1.9% 0.6% 5.1%
Academia 0.1% 0 0.6% 0 0 0 0 0 0 0
Media 0.3% 0 7.3% 0.5% 0 0 0 0.2% 0 0
Auditors and Oversight Boards 0 0 0.8% 0 0 0 0 0 0 0
Advocacy Groups 1.5% 0 1.0% 4.2% 0 0.1% 0 0 0 0
Table 2: Stakeholder involvement across reparative actions. AI corporations mainly engaged in public communication, with limited involvement in structural remedies. Affected communities participated infrequently, and third-party actors helped surface harm but rarely shaped resolution. For full action names, see Table 1.
@@ Discussion
In this section, we first use justice frameworks to assess the reparative actions in the taxonomy, examining how efforts toward remedy and reform align with punitive, restorative, and transformative ideals. Next, we examine which stakeholders engage in the reparation process and how power is distributed; in particular, we highlight the constrained role of affected communities and the critical, but often overlooked, contributions of civil society actors. Across both areas, we identify key gaps and offer recommendations for more inclusive and effective approaches to reparation.
@@ A Vacuum of Justice: Evaluating Reparative Actions Through Justice Frameworks
While we categorize reparative actions by justice-oriented goals, this framing also helps us understand how these ideals are realized in practice. Justice frameworks are not prescriptive but serve as interpretive tools to identify shortcomings and possibilities for more meaningful redress. This approach aligns with calls in Responsible AI to develop context-sensitive, relational models of accountability (Metcalf et al. 2023). Here, we examine two dimensions: how actions under Remedy reflect principles of punitive and restorative justice, and how those under Reform correspond to transformative justice.
Remedy for Past Harm: Punishment and Repair
@@
Corrective justice traditions emphasize that those responsible for harm should be held accountable and that concrete efforts should be made to repair the losses experienced by those affected (Feinberg 2019; Zehr 2015). These principles underlie both punitive and restorative justice, which guided our analysis of actions under the goal of Remedy.
@@
Punitive justice holds that wrongdoers should face consequences proportionate to the harm they caused (Feinberg 2019). In our dataset, however, punitive responses were rare. Only 10% of cases involved fines, bans, or dismissals, typically enforced by regulators or courts. We found much less evidence of internal accountability, such as companies publicly disclosing disciplinary action or self-imposed sanctions. While internal measures may occur, the lack of transparency makes them difficult to assess. This aligns with prior research showing that AI professionals often view accountability as diffuse and located outside their individual or organizational control (Lancaster et al. 2024). Moreover, when punitive actions take place, they generally relied on traditional legal frameworks, such as privacy or copyright violations, rather than mechanisms tailored to AI-specific harms. This limits the effectiveness of punitive measures in addressing the unique and evolving risks of AI systems.
@@
Restorative justice focuses on repairing harm by centering the needs and agency of those affected and by enabling acknowledgment, dialogue, and restitution (Zehr 2015). Yet restorative practices were similarly scarce. Only 7% of cases involved direct remedy to affected communities in the form of compensation. Few included meaningful input from those harmed in shaping the response. While public apologies and statements of intent were more common, they were often vague, symbolic, and lacked follow-through. As prior work has noted, such gestures may function more as reputational risk management than as genuine accountability (Metcalf, Moss et al. 2019; Green 2021). These patterns reflect a broader gap in current practices: even when harm is acknowledged, efforts to rebuild trust or restore relationships are limited. These patterns point to a deeper issue: even when harm is recognized, there is little investment in restoring relationships or rebuilding trust. Despite growing calls for participatory, victim-centered approaches (Ajmani et al. 2024; Ganesh and Moss 2022; Velkova and Kaun 2021), most responses failed to return agency to affected communities, leaving the core aims of restorative justice unmet.
Reform for Structural Harm: Systemic Changes
@@
Actions under the goal of Reform targeted institutional or systemic conditions, such as organizational policies or algorithm designs. These responses align with the aims of transformative justice, which seeks to prevent future harm by confronting and altering the structural and social conditions that allow harm to persist (Morris 2000).
@@
In our dataset, 22% of cases involved actions that could be interpreted as targeting structural factors. These included product recalls, system redesigns, and law and policy change. However, many of these interventions were narrow in scope. They rarely addressed upstream drivers such as exploitative data practices, profit incentives that reward harmful behavior, or recurring harms across platforms. Most changes were confined to a single organization and lacked coordination across the broader ecosystem. In some cases, structural change took the form of product withdrawal rather than substantive redesign—what has been critiqued as a superficial fix that evades accountability (Johnson et al. 2024). These patterns echo broader critiques that ethics initiatives in the tech industry often serve to preempt regulation and manage public image rather than drive systemic accountability (Metcalf, Moss et al. 2019; Green 2021).
Remedy and Reform: Reparation of future vs reparation of the past
@@
A key distinction in our taxonomy is between reparative actions that address past harm (Remedy) and those that seek to restructure systems to reduce recurrence (Reform). This distinction helps clarify a gap noted in existing scholarship: systemic changes cannot replace direct redress to those harmed. Recent critiques have pointed out that widely used accountability practices—such as auditing and red teaming—often serve to mitigate risk rather than ensure meaningful reparation (Gillespie et al. 2024; Feffer et al. 2024). Similarly, AI policy frameworks frequently prioritize transparency and procedural oversight while lacking enforceable requirements for remediation (Wright et al. 2024). Centering reparation requires that both individual and systemic dimensions be addressed. Remedy and reform must work together to ensure that harms are not only acknowledged but actively repaired through inclusive and sustained action.
Uneven Participation in Reparation: Role of Affected Community and Third Parties
Our taxonomy reveals uneven stakeholder participation across the reparation process. Companies often engage early through public communication but are less involved in providing remedy or pursuing reform. Affected communities help surface harm and assign responsibility but are rarely included in shaping responses. Civil society and the public, though not directly harmed, play crucial roles in defining accountability. This uneven distribution of responsibility reflects broader gaps in the design of AI accountability infrastructures, where institutional responses often overlook the contributions of affected individuals and third-party actors (DeVrio, Eslami, and Holstein 2024).
@@
Moreover, it is difficult for affected individuals to mobilize collectively, particularly when harms are diffuse, individualized, or legally ambiguous. Lawsuits, one of the few formal channels for redress, were more often initiated by regulators or advocacy groups than by those directly affected. This reflects deeper structural barriers: affected communities often lack the resources, visibility, or institutional support to demand meaningful accountability. Even when users engage in public contestation of AI harm, these efforts rarely connect with formal mechanisms of redress (Velkova and Kaun 2021; Ganesh and Moss 2022). Addressing this gap requires rethinking reparation as a participatory process that centers the agency and priorities of those harmed. Without such inclusion, reparation risks reinforcing the very power asymmetries it aims to challenge.
@@
Our analysis highlights a broader ecosystem of actors who influence accountability, especially where formal reparation mechanisms are lacking. While existing literature on multi-stakeholder AI governance often focuses on developers and directly affected users (Morley et al. 2020; Rakova et al. 2020; Mohamed, Png, and Isaac 2020), we extend this view to include third-party actors such as advocacy groups, journalists, academics, and members of the public. Although not involved in AI development, these actors play key roles in initiating, sustaining, and shaping responses to harm through public oversight and intervention.
Instead of being passive observers, they frequently trigger the recognition of harm and apply pressure that compels institutional action. Advocacy groups raise awareness and mobilize campaigns (Contreras 2023); journalists conduct investigations and frame incidents as systemic (Angwin et al. 2016); academics produce audits and critiques (Deng et al. 2025; Shen et al. 2021; Cabrera et al. 2021); and the public amplifies these efforts through petitions and media engagement (DeVrio, Eslami, and Holstein 2024). These roles echo “outsider oversight” models (Raji et al. 2022) and mirror civil society’s influence in other domains (Lucaj, Van Der Smagt, and Benbouzid 2023; Gillespie et al. 2024).
Despite their importance, third-party efforts often operate in isolation from institutional processes. These actors may surface harm or apply public pressure but are rarely connected to formal mechanisms such as legal action, compensation, or policy reform. Prior work has noted that civil society interventions often remain fragmented and reactive, rather than integrated into sustained oversight (Raji et al. 2022; Lucaj, Van Der Smagt, and Benbouzid 2023; Gillespie et al. 2024). This disconnect partly reflects structural barriers, as these actors are not always directly affected and lack formal roles in governance. Yet their involvement is essential, especially given the wider societal impact of AI harms. To bridge this gap, institutions should create formal mechanisms for collaboration such as supporting independent audits, consulting civil society in policy design, and enabling their sustained participation in reparative processes (Lam et al. 2022; Feffer et al. 2024). Better integration can turn public pressure into institutional change and ensure reparation aligns with long-term public interest.
@@ Conclusion
This paper introduces a taxonomy of AI harm reparation and applies it to a large-scale analysis of post-incident responses. The taxonomy structures reparative actions into four stages, providing a lens to examine how harm is acknowledged, addressed, remedied, and potentially transformed over time. Our analysis highlights not only the uneven distribution of reparative efforts but also the limitations of current responses in delivering meaningful redress. Drawing on theories of corrective justice, we argue that most existing responses fall short of punitive, restorative, or transformative ideals. Rather than treating reparation as a discrete or one-time intervention, we call for a more sustained and systemic approach. This includes designing mechanisms that support affected communities, hold powerful actors accountable, and ensure accountability in carrying out law and policy.
@@ Appendix B Appendix B: Stakeholder Definitions
Stakeholder Definitions
@@
Table 3: Primary stakeholder categories referenced in our taxonomy.

### v2

@@ Acknowledgment
Advocacy groups frequently pursue legal accountability when other accountability channels fail. For instance, the UK-based Public Law Project issued a formal pre-action letter challenging the Home Office’s algorithm for flagging “sham marriages,” citing discrimination and GDPR violations. As of the latest reporting, the case had not reached a resolution (A1389).
Affected Individuals, including public figures and professionals, also initiated legal actions, typically over severe harm or personal rights violations. For example, several religious authors sued Meta and Microsoft for allegedly generating AI content derived from their copyrighted works (A1149). In another case, Megan Garcia filed a negligence lawsuit against Character AI after her son died by suicide following emotional attachment to a chatbot (A1781). Actress Scarlett Johansson also sued an AI developer for unauthorized use of her likeness (A1165).
@@
While legal actions provide a formal avenue for redress, our findings show they are typically slow-moving, legally complex, and more accessible to well-resourced individuals or organizations. Cases initiated by regulators are far more likely to result in concrete consequences than those led by private actors.
@@
Removal of Content or Product Recall
@@
While our taxonomy identifies which stakeholder groups initiate each type of reparative action, it is organized around actions and does not reveal how individual stakeholders contribute across the broader set of actions. To complement this, Table 2 presents the frequency with which each stakeholder group initiates each action, with percentages reflecting their occurrence across all 1,060 incidents. This comparative view highlights patterns of responsibility and exposes disparities in who drives different forms of reparation.
Stakeholder Group Outcry Comm. Invest. Legal. Repr. Comp. Recall Design Discont. Policy
Perpetrators – AI Corporations 46.1% 6.8% 0.9% 1.1% 5.0% 9.2% 10.3% 3.7% 2.7%
Perpetrators – End-Users 0.1% 7.2% 0.7% 0.4% 0.3% 0.9% 1.5% 0.1% 0.5% 0.4%
Affected Users 2.6% 0.1% 7.9% 0.1%
Affected Non-Users 0.6% 2.6%
General Public 6.6%
Regulators and Government Bodies 0.2% 0.3% 20.8% 2.2% 9.5% 1.2% 1.9% 0.8% 0.6% 5.1%
Academia 0.1% 0.6%
Media 0.3% 7.3% 0.5% 0.2%
Auditors and Oversight Boards 0.8%
Advocacy Groups 1.5% 1.0% 4.2% 0.1%
Table 2: Stakeholder involvement across reparative actions. Percentages represent the proportion of all 1,060 incidents. Dashes indicate no involvement; bolded values show each stakeholder’s highest level of engagement. AI corporations focused on communication, while structural remedies saw limited participation. Affected communities engaged infrequently; third parties surfaced harm but rarely shaped outcomes. Please refer to Table 1 for complete action names.
@@ Discussion
This section examines how post-incident reparation efforts reflect, or fall short of, the pursuit of justice in AI harm. We begin by analyzing how actions aimed at remedy and reform correspond to theories of punitive, restorative, and transformative justice. While many of these actions seek to address harm, they often lack clear accountability, provide limited reparation to affected communities, or fail to engage with the structural conditions that produced the harm. We then turn to participation in the reparation process, focusing on how affected communities and third-party actors contribute to reparation, though often through informal or constrained roles.
@@ A Vacuum of Justice: Evaluating Reparative Actions Through Justice Frameworks
Justice frameworks not only help categorize reparative actions but also reveal how current responses fall short of meaningful accountability and redress. Rather than offering prescriptions, these frameworks serve as interpretive tools for identifying limitations and possibilities. This approach aligns with calls in Responsible AI to develop context-sensitive, relational models of accountability (metcalf2023taking). Below, we examine how actions under Remedy reflect principles of punitive and restorative justice, and how those under Reform correspond to transformative justice.
Remedy for Past Harm: Rare Accountability and Exclusion of the Harmed
@@
Corrective justice traditions emphasize that those responsible for harm should be held accountable and that concrete efforts should be made to repair the losses experienced by those affected (feinberg2019expressive; zehr_little_2015). These principles underlie both punitive and restorative justice, which guided our analysis of actions under the goal of Remedy.
@@
Punitive justice holds that wrongdoers should face consequences proportionate to the harm they caused (feinberg2019expressive). In our dataset, however, punitive responses were rare. Only 10% of cases involved fines, bans, or dismissals, typically enforced by regulators or courts. We found much less evidence of internal accountability, such as companies publicly disclosing disciplinary action or self-imposed sanctions. While internal measures may occur, the lack of transparency makes them difficult to assess. This aligns with prior research showing that AI professionals often view accountability as diffuse and located outside their individual or organizational control (lancaster2024s). Moreover, when punitive actions take place, they generally relied on traditional legal frameworks, such as privacy or copyright violations, rather than mechanisms tailored to AI-specific harms. This limits the effectiveness of punitive measures in addressing the unique and evolving risks of AI systems.
@@
Restorative justice focuses on repairing harm by centering the needs and agency of those affected and by enabling acknowledgment, dialogue, and restitution (zehr_little_2015). Yet restorative practices were similarly scarce. Only 7% of cases involved direct remedy to affected communities in the form of compensation. Few included meaningful input from those harmed in shaping the response. While public apologies and statements of intent were more common, they were often vague, symbolic, and lacked follow-through. As prior work has noted, such gestures may function more as reputational risk management than as genuine accountability (metcalf2019owning; green2021contestation). These patterns reflect a broader gap in current practices: even when harm is acknowledged, efforts to rebuild trust or restore relationships are limited. Despite growing calls for participatory, victim-centered approaches (ajmani2024data; ganesh2022resistance; velkova_algorithmic_2021), most responses failed to return agency to affected communities, leaving the core aims of restorative justice unmet.
Reform for Structural Harm: Structural Responses Without Structural Change
@@
Actions under the goal of Reform aimed to address institutional or systemic conditions, such as organizational policies or algorithm designs. These responses align with the aims of transformative justice, which seeks to prevent future harm by confronting and altering the structural and social conditions that allow harm to persist (morris2000stories). However, in practice, such actions often fell short of these goals.
@@
In our dataset, 20% of cases involved actions that could be interpreted as targeting structural factors. These included product and feature discontinuation, system redesigns, and law and policy change. However, many of these interventions were narrow in scope. They rarely addressed upstream drivers such as exploitative data practices, profit incentives that reward harmful behavior, or recurring harms across platforms. Most changes were confined to a single organization and lacked coordination across the broader ecosystem. In some cases, structural change took the form of product discontinuation rather than substantive redesign—what has been critiqued as a superficial fix that evades accountability (johnson_fall_2024). These patterns echo broader critiques that ethics initiatives in the tech industry often serve to preempt regulation and manage public image rather than drive systemic accountability (metcalf2019owning; green2021contestation).
Remedy and Reform as Complementary Forms of Reparation
@@
A key distinction in our taxonomy is between reparative actions that address past harm (Remedy) and those that restructure systems to reduce recurrence (Reform). This distinction helps clarify a gap noted in existing scholarship: systemic changes cannot replace direct redress to those harmed. Recent critiques have pointed out that widely used accountability practices—such as auditing and red teaming—often mitigate risk rather than ensure meaningful reparation (gillespie2024ai; feffer_red-teaming_2024). Similarly, AI policy frameworks frequently prioritize transparency and procedural oversight while lacking enforceable requirements for remediation (wright2024null). Centering reparation requires that both individual and systemic dimensions be addressed. Remedy and reform must work together to ensure harms are not only acknowledged but actively repaired through inclusive and sustained action.
Who Gets to Repair? Participation in AI Reparation
Beyond mapping past responses, our taxonomy offers a forward-looking tool for guiding more accountable and participatory approaches to AI harm reparation. For affected communities, it provides language to articulate what they are owed and where redress is lacking. For companies and policymakers, it highlights neglected levers for remedy and reform. Researchers and auditors may use it to assess the distribution of responsibility, benchmark institutional responses, and track progress over time.
Yet while our taxonomy outlines a range of potential contributions, carrying out these roles in practice often faces structural and institutional barriers, as seen across our AI incident analysis. These challenges limit the ability of stakeholders, particularly those not involved in AI development or regulatory processes, to meaningfully engage in reparation. Below, we highlight two groups often marginalized in reparative processes: affected communities and third-party actors.
@@
Moreover, it is difficult for affected individuals to mobilize collectively, particularly when harms are diffuse, individualized, or legally ambiguous. Legal actions, one of the few formal channels for redress, were more often initiated by regulators or advocacy groups than by those directly affected. This reflects deeper structural barriers: affected communities often lack the resources, visibility, or institutional support to demand meaningful accountability. Even when users engage in public contestation of AI harm, these efforts rarely connect with formal mechanisms of redress (velkova_algorithmic_2021; ganesh2022resistance). Addressing this gap requires rethinking reparation as a participatory process that centers the agency and priorities of those harmed. Without such inclusion, reparation risks reinforcing the very power asymmetries it aims to challenge.
@@
Our analysis highlights a broader ecosystem of actors who influence reparation and accountability, especially where formal mechanisms are lacking. While multi-stakeholder AI governance research often centers on developers and directly affected users (morley2020initial; rakova_where_2020; mohamed2020decolonial), we extend this view to include civil society and the public, including advocacy groups, journalists, academics, and community members.
Prior Responsible AI scholarship has documented what these actors do to resist or prevent harm, such as advocacy groups mobilizing campaigns (contreras2023civilrightsai), journalists framing incidents as systemic issues (angwin2016machine), academics producing audits and critiques (deng2025weaudit; shen_everyday_2021; cabrera2021discovering), and the public amplifying these efforts through petitions and media engagement (devrio_building_2024). Building on this work, our contribution is to place these activities within the reparation process and show how they operate as oversight and public pressure that can create pathways toward remedy and reform.
Yet despite their impact, these interventions remain loosely connected to formal mechanisms such as legal remedies, compensation, or policy reform. Prior work notes that civil society’s role is often fragmented and reactive (raji2022outsider; lucaj2023ai; gillespie2024ai), reflecting structural barriers: these actors are not directly affected by every incident and lack formal standing in governance. Still, their participation is essential given the societal reach of AI harms. Institutions can bridge this gap by creating channels for collaboration—supporting independent audits, consulting civil society in policy design, and enabling sustained involvement in reparative processes (lam_end-user_2022; feffer_red-teaming_2024). Better integration can transform public pressure into institutional change and align reparation with the broader public interest.
@@ Conclusion
This paper introduces a taxonomy of AI harm reparative actions and applies it to a large-scale analysis of post-incident responses. The taxonomy traces how harm is acknowledged, addressed, remedied, and, in some cases, transformed over time. Our findings reveal the uneven distribution of reparative efforts and the limits of current responses in achieving meaningful redress. We argue for more systemic approaches that center affected communities, integrate third-party oversight, and tie commitments to enforceable accountability.
@@ Appendix B Appendix B: Stakeholder Definitions
Table 3: Definitions of primary stakeholder categories referenced in our taxonomy.